Used Cars - Actual Data: Below are the scatterplots, regression equations, and corresponding statistics for mileage, model year, and price for 15 different Honda Civics found on craigslist in May 2012.

         

Mileage -vs- Price:
x = Mileage (in thousands of miles)
y = Price (in dollars)

correlation coefficient:
r = −0.886

regression equation:
ŷ = −76.12x + 15,585.8

sample size:
n = 15

         

Model Year -vs- Price:
x = Model Year
y = Price (in dollars)

correlation coefficient:
r = 0.899

regression equation:
ŷ = 992.43x − 1,980,582.4

sample size:
n = 15

Suppose you see a 2004 Honda Civic with 160 thousand miles on it. Estimate a reasonable price for this car via the following methods.

(a) Estimate the price using the mileage. Round your answer to the nearest dollar.
$

(b) Estimate the price using the year. Round your answer to the nearest dollar.
$

(c) Estimate the price using the multiple linear regression equation given by

ŷ = 717.3x1 − 42.7x2 − 1,424,349

where x1 is the model year and x2 is the mileage (in thousands). Round your answer to the nearest dollar.
$

(d) Which of the following statements are valid?

The estimate based only on mileage (part a) is too low because it doesn't consider the model year.The estimate based only on the year (part b) is too high because it doesn't consider the mileage.     The estimate from part (c) considers both variables (year and mileage) to produce a better estimate.All of these are valid statements.
Answer #1

a)

Estimated price =-76.12*160+15585.8 = 3407

b)

Estimated price =992.43*2004-1980582.4 =8247

c)

Estimated price =717.3*2004-42.7*160-1424349 =6288

d)

All of these are valid statements.